OpenStudy (anonymous):

The expression 3r - 6 can be written as _____. A. 3(r- 2) B. 3(r- 6) C. 3(3r- 6)

OpenStudy (anonymous):

3r-6=3*r-(2*3)=3*(r-2) so A

OpenStudy (anonymous):

thx ur getting best anwser

OpenStudy (anonymous):

gallent can you help me on another problem

OpenStudy (anonymous):

A